OpenStudy (anonymous):

What is the slope of the line 12x + 4y = 24 What is the answer?

OpenStudy (anonymous):

y=-3x+6

OpenStudy (anonymous):

the slope is -3

OpenStudy (haseeb96):

this is the formula of slope y=mx+b 12x+4y=24 separate the y 4y=24-12x 4y=4(6-3x) since;..... 4\4=1 y=6-3x y=-3x+6 m=-3 this is the slope

OpenStudy (anonymous):

Put it into y=mx+b form, 12x + 4y = 24 then move the 12x to the other side to put it in slope intercept form, which would make it -12x 4y = -12x + 24 then divide each number by 4 to get Y by itself y= -3x+6 -3 is the slope (m in the y=mx+b formula)